physics lab A device is created in the lab which measures the
Answer #1

question 1) you have to convert the mass to kg

p_A = (0.2533kg)(0.87m/s) = 0.22 kg.m/s

Combined uncertainty :

u = \sqrt{u_m^2 + u_v^2} = \sqrt{(0.05\times10^{-3})^2 + (0.02)^2} = 0.02

question 2) same procedure:

p_B = m_B v_B = (0.2467kg)(0.51m/s) = 0.13kg.m/s

and the uncertainty is the same.

Question 3) Conservation of linear momentum

p_{A,i} + p_{B,i} = p_f

m_A v_{A,i} + m_B v_{B,i} = (m_A+m_B)v_f

0.22kg.m/s - 0.13kg.m/s= (0.2533kg+0.246.7kg)v_f

Momentum for B is negative , because it's moving to the left.

Solving for final velocity you get:

v_f = 0.18m/s

The combined uncertainty is for this expression:

v_f = \frac{p_A+p_B}{m_A+m_B}

The uncertainty for the numerator is:

u_1 = \sqrt{u_{p_A}^2 + u_{p_B}^2} = \sqrt{(0.02)^2+(0.02)^2} = 0.028

and for the denominator:

u_2 = \sqrt{u_{m_A}^2 + u_{m_B}^2} = \sqrt{(0.05\times10^{-3})^2+(0.05\times10^{-3})^2} = 7.1\times10^{-5}

The uncertainty for the quotient is:

u = \sqrt{u_{1}^2 + u_{2}^2} = \sqrt{(0.028)^2+(7.1\times10^{-5})^2} = 0.028